OpenStudy (anonymous):

How do I make y = 2-(4/(X+1)) be in the form x= Please help I got to y-2 = -(4/X+1)

OpenStudy (ribhu):

4/(x+1) = 2-y

OpenStudy (ribhu):

(x+1)/4 = 1/2-y

OpenStudy (ribhu):

x+1 = 4/2-y

OpenStudy (ribhu):

x = (4/2-y)-1

OpenStudy (ribhu):

x = (4-2+y)/2-y

OpenStudy (anonymous):

Stop.

OpenStudy (ribhu):

x= (2+y)/(2-y)

OpenStudy (ribhu):

i solved it for u @OpiGeode

OpenStudy (ribhu):

why stop? @OpiGeode u reply